# Intentional Gesture: Deliver Your Intentions with Gestures for Speech

### v1

@@ 1 Speaker All Speakers Novel 5 speakers
CaMN [14] $0.604$ $0.676$ $9.97$ $0.512$ $0.200$ $5.58$ $0.812$ $0.563$ $6.71$
EMAGE [13] $0.570$ $0.793$ $11.41$ $0.692$ $0.284$ $6.06$ $0.936$ $0.643$ $7.47$
Audio2Photoreal [22] $1.02$ $0.550$ 12.47 $0.849$ $0.326$ $6.24$ $1.012$ $0.464$ $5.97$
RAG-Gesture [21] $0.879$ $0.730$ $12.62$ $0.447$ 0.471 $9.03$
GestureLSM [15] $0.408$ $0.714$ $13.24$ $0.446$ $0.525$ $9.23$ $0.664$ $0.621$ $10.45$
@@ 5.2 Qualitative Comparisons
Evaluation Results As depicted in Figure 6, our approach generates gestures that exhibit improved rhythmic alignment and a more natural appearance compared to existing methods. For example, when conveying the phrase “helpful for me”, our method directs the subject to extend left hand forward, effectively representing the intention of an explanation. In contrast, competing methods fail to capture this nuance, often generating static or unnatural poses where one or both arms remain down.
@@
User Study. We conducted a user study with 20 participants and 160 video samples, 40 from each of GestureLSM [15], EMAGE [13], CAMN [14], and Ours. Each participant viewed the videos in a randomized order and rated them on a scale of 1 (lowest) to 5 (highest) based on three criteria: (1) realness, (2) speech-gesture synchrony, and (3) smoothness. As shown in Tab.2, Our method outperforms other methods across all criteria, achieving higher Mean Opinion Scores (MOS).
@@ 6 Ablation Studies and Analysis
We explore the hierarchical audio-intention and motion alignment, the design of tokenizer and generator in this section. We defer the further analysis of the quality of annotation effects on learning, model generalization ability, sequence length effect, and quantizer analysis to Appendix E.
Figure 7: Hierarchical Alignment Analysis. Retrieval-based evaluation of H-AuMoCLIP across low-level (rhythmic) and high-level (semantic) features. We report Recall@1/10 under per-batch and global settings for both known and unknown speakers. Incorporating intentional semantics improves retrieval performance, especially for high-level alignment and with few training speakers. Increasing speaker diversity further enhances generalization in high-level retrieval.
@@ 6.1 Evaluation of Hierarchical AuMoCLIP
We evaluate H-AuMoCLIP using retrieval-based metrics that measure how well audio features align with corresponding motion features at both low and high levels. We report Recall@1 and 10 under two settings: per-batch (within batch of 128) and global (across the entire test set).
Low-level retrieval evaluates frame-wise alignment by computing cosine similarity between audio and motion features at each timestep, within a local temporal window $(i\pm 16)$. High-level retrieval evaluates sequence-level alignment using mean-pooled audio and motion features for full-clip matching. Please see more details in TANGO [12]. We evaluate generalization to both seen (in-domain) and unseen (out-of-domain) speakers, assessing robustness across speaker identities.
@@
Speaker Diversity Enhances Generalization. Training on multiple speakers significantly improves generalization to unseen speakers. Even with the same total training data, models trained on more speaker identities achieve better high-level retrieval, suggesting that speaker diversity encourages learning more robust gesture-speech mappings, as shown in Appendix E.
Low-Level Alignment Remains Stable. In contrast, low-level beat synchronization metrics show minimal gains from increased data or speaker diversity, indicating that rhythm alignment depends more on prosodic regularities than on inter-speaker variation.
Figure 8: Tokenizer ablation. We perform both global and local grid searches to study the effects of codebook design choices on rFGD (×$\times 10^{3}$). (a)–(c): Global sweeps varying one factor at a time; (d): Local grid search over code count and dimension. All results confirm consistent trends: 8 codebooks, a code dimension of 32, and 8192 codes yield optimal or near-optimal performance.

### v2

@@ 1 Speaker All Speakers Novel 5 speakers
CaMN Liu et al. (2022b) $0.604$ $0.676$ $9.97$ $0.512$ $0.200$ $5.58$ $0.812$ $0.563$ $6.71$
EMAGE Liu et al. (2023) $0.570$ $0.793$ $11.41$ $0.692$ $0.284$ $6.06$ $0.936$ $0.643$ $7.47$
Audio2Photoreal Ng et al. (2024) $1.02$ $0.550$ 12.47 $0.849$ $0.326$ $6.24$ $1.012$ $0.464$ $5.97$
RAG-Gesture Mughal et al. (2025) $0.879$ $0.730$ $12.62$ $0.447$ 0.471 $9.03$ - - -
GestureLSM Liu et al. (2025a) $0.408$ $0.714$ $13.24$ $0.446$ $0.525$ $9.23$ $0.664$ $0.621$ $10.45$
@@ 5.2 Qualitative Comparisons
Evaluation Results As depicted in Fig. 6, our approach generates gestures that exhibit improved rhythmic alignment and a more natural appearance compared to existing methods. For example, when conveying “helpful for me”, our method directs the subject to extend left hand forward, effectively representing the intention of an explanation. In contrast, competing methods fail to capture this nuance, often generating static or unnatural poses where one or both arms remain down.
@@
User Study. We conducted a user study with 20 participants and 160 video samples, 40 from each of GestureLSM Liu et al. (2025a), EMAGE Liu et al. (2023), CAMN Liu et al. (2022b), and Ours. Each participant viewed the videos in a randomized order and rated them on a scale of 1 (lowest) to 5 (highest) based on three criteria: (1) realness, (2) speech-gesture synchrony, and (3) smoothness. As shown in Tab.2, Our method outperforms other methods across all criteria, achieving higher Mean Opinion Scores (MOS).
Figure 7: Hierarchical Alignment Analysis. Retrieval-based evaluation of H-AuMoCLIP across semantic features. We report Recall@1/10 under per-batch and global settings for both known and unknown speakers. Incorporating intentional semantics improves retrieval performance with few training speakers. Increasing speaker diversity further enhances generalization for retrieval.
@@ 6 Ablation Studies and Analysis
We explore the hierarchical audio-intention and motion alignment, the design of tokenizer and generator in this section. We defer the further analysis of the quality of annotation effects on learning, model generalization ability, sequence length effect, and quantizer analysis to Appendix.
@@ 6.1 Evaluation of Hierarchical AuMoCLIP
We evaluate H-AuMoCLIP using retrieval-based metrics that measure how well audio and intentional features align with corresponding motion features. We report Recall@1 and 10 under two settings: per-batch (within batch of 128) and global (across the entire test set). The retrieval evaluates sequence-level alignment using mean-pooled audio and motion features for full-clip matching. We evaluate generalization to both seen (in-domain) and unseen (out-of-domain) speakers, assessing robustness across speaker identities.
@@
Speaker Diversity Enhances Generalization. Training on multiple speakers significantly improves generalization to unseen speakers. Even with the same total training data, models trained on more speaker identities achieve better high-level retrieval, suggesting that speaker diversity encourages learning more robust gesture-speech mappings, as shown in Appendix.